# EVQAScore: Efficient Video Question Answering Data Evaluation

### v1

@@ 1 Introduction
To address these challenges, we introduce EVQAScore, which utilizes LLMs to extract keywords for enhanced video captioning and video QA data evaluation. To support the evaluation of longer videos, we efficiently compute EVQAScore through uniform frame sampling, which reduces computational costs by a factor of 30 without affecting results. Compared to previous methods such as EMScore [34] and PAC-S [33], our approach can be applied to video QA evaluation and achieves superior visual understanding through keyword extraction, outperforming the TF-IDF technique used by EMScore and PAC-S. Additionally, due to its efficiency and keyword extraction capabilities, our method can handle extremely long captions and videos, a feat not possible for EMScore and PAC-S because of the 77-word limit imposed by CLIP.
The core contributions of this paper are summarized as follows:
  • New Perspective: We introduce a novel approach, EVQAScore, designed for the evaluation of Video QA and video caption data quality. To the best of our knowledge, this is the first systematic effort to evaluate the quality of Video QA data.
  • New Method: Our approach leverages LLMs and keyword extraction techniques to improve the understanding of QA data in Video QA evaluation. Additionally, we employ uniform frame sampling to enhance scalability, allowing EVQAScore to be applied to extremely long videos and large datasets.
@@
  • SOTA Performance for Video Caption Evaluation: We evaluated the performance of our method on the VATEX-Eval benchmark for Video Caption data, achieving a Kendall correlation of 32.8 and a Spearman correlation of 42.3, which are 4.7 and 5.9 points higher, respectively, than the previous method PAC-S++.
  • SOTA Performance for Video QA Evaluation: We constructed a dataset comprising 400K video entries, consisting of 200K noisy samples and 200K high-quality samples. Using our EVQAScore and the previous SOTA model PAC-S, we filtered the data, retrieving a significantly lower amount of noisy data than PAC-S. Additionally, when evaluated on ActivityNet, MSRVTT, MSVD, TGIF, MVBench, and two longer benchmarks—VideoChatGPT Bench and VideoChatGPT Bench Diverse—our EVQAScore outperformed PAC-S and 100% of data with only 12.5% of data.
@@ 3 Method
Figure 2: Illustration of the computation of EMScore which utilize coarse to fine video evaluation. We utilize keyword extraction and then use frames and keywords as ground truth for
@@ 3.2 Efficiency of EVQAScore
In our experiment we select sampling interval $l$ to be 10, 20, and 30. We present the result in Figure
@@ 3.3 Video caption evaluation
where $f_{V}$ is the embedding of the video, which is calculated by pooling the feature of video frames and caption, respectively. $f_{X}$ is the feature of caption or QA pair.
@@
$P(V,X)_{f}=P\left(V,K\right)_{f}=\frac{1}{|X|}\sum_{k_{j}\in K}\max_{v_{i}\in V}f_{v_{i}}^{\top}f_{k_{j}}$
$R(V,X)_{f}=R\left(V,K\right)_{f}=\frac{1}{|V|}\sum_{v_{i}\in V}\max_{k_{j}\in K}f_{v_{i}}^{\top}f_{k_{j}}$
@@
$\text{EVQAScore}(X,V)_{f}=2\frac{P(V,X)_{f}\cdot R(V,X)_{f}}{P(V,X)_{f}+R(V,X)_{f}}$
@@ 4.1 Experimental Setting
We employed the positive-augmented CLIP visual encoder [33] to extract embeddings from individual frames, applying average pooling to represent the entire video. Similarly, we used the corresponding textual encoder to generate token embeddings, producing a comprehensive caption representation. Additionally, we utilized the LLaMA3.1 8B-Instruct model to identify key terms from question-answer pairs, highlighting the essential elements of the dialogue.
@@
We used EMScore, PAC-S, and PAC-S++ as baselines. Both PAC-S and PAC-S++ do not extract keywords or frames, while PAC-Score leverages the positive-augmented CLIP model to enhance representation.
@@
For LLaMA3, we applied the official hyperparameters from the repository. For the CLIP model, we used the clip-ViT-B-32 variant for keyframe selection. All experiments were conducted on a machine equipped with 8 NVIDIA H800 GPUs, a 192-core CPU, and 960GB of memory.
@@ 4.4 Video QA Data Filtering
Table 7: Performance of VideoLLaVA and VideoLLaMA2 on MVBench using filtered by PAC-S and EVQAScore.
@@ 4.6 Case Study: Evaluating Video Captions
We summarize the first case in Figure 1. Given the question: ”What is the man doing in the video?” the first answer contains only one word and is of low quality, while the second answer is longer and of higher quality. Compared with CLIPScore (using the pooling feature of video), EMScore, PAC-S, and EVQAScore, the video QA pair indicates that our method is the only approach capable of recognizing the second answer as being of higher quality. This demonstrates the robustness of our method.

### v2

@@ 1 Introduction
To address these challenges, we introduce EVQAScore, which utilizes LLMs to extract keywords for enhanced video captioning and video QA data evaluation. To support the evaluation of longer videos, we efficiently compute EVQAScore through uniform frame sampling, which reduces computational costs by a factor of 30 without affecting results. Compared to previous methods such as EMScore [34] and PAC-S [33], our approach can be applied to video QA evaluation and achieves superior visual understanding through keyword extraction, outperforming the TF-IDF technique used by EMScore and PAC-S. Additionally, due to its efficiency and keyword extraction capabilities, our method can handle extremely long captions and videos, a feat not possible for EMScore and PAC-S because of the 77-word limit imposed by CLIP. The core contributions of this paper are summarized as follows:
  • New Method: We introduce a novel approach, EVQAScore, designed for the evaluation of Video QA and video caption data quality. To the best of our knowledge, this is the first systematic effort to evaluate the quality of Video QA data. Our approach leverages LLMs and keyword extraction techniques to improve the understanding of QA data in Video QA evaluation. Additionally, we employ uniform frame sampling to enhance efficiency and scalability, allowing EVQAScore to be applied to extremely long videos and large datasets.
@@
  • SOTA Performance for Video Caption Evaluation: We utilize LLM for keyword extraction for better caption and QA understanding. We evaluated the performance of our method on the VATEX-Eval benchmark for Video Caption data, achieving a Kendall correlation of 32.8 and a Spearman correlation of 42.3, which are 4.7 and 5.9 points higher than the previous method PAC-S++.
  • SOTA Performance for Video QA Evaluation: We constructed a dataset comprising 400K video entries, consisting of 200K noisy samples and 200K high-quality samples. Using our EVQAScore and the previous SOTA model PAC-S, we filtered the data, retrieving a significantly lower amount of noisy data than PAC-S. As shown in Figure 1, our method is the only method that can identify low-quality video QA data. Additionally, when evaluated on ActivityNet, MSRVTT, MSVD, TGIF, MVBench, and two longer benchmarks—VideoChatGPT Bench and VideoChatGPT Bench Diverse—our EVQAScore outperformed PAC-S and 100% of data with only 12.5% of data.
@@ 3 Method
Figure 2: Illustration of the computation of EMScore which utilize coarse to fine video evaluation. We utilize keyword extraction and then use frames and keywords as ground truth for precision and recall calculation.
@@ 3.2 Efficiency of EVQAScore
In our experiment we select sampling interval $l$ to be 10, 20, and 30. We present the result in Table 1.
@@ 3.3 Video caption evaluation
where $f_{V}$ is the embedding of the video, which is calculated by pooling the feature of video frames in sampled frames $S$, respectively. $f_{X}$ is the feature of caption or QA pair.
@@
$P(V,X)_{f}=P\left(S,K\right)_{f}=\frac{1}{|X|}\sum_{k_{j}\in K}\max_{v_{i}\in S}f_{v_{i}}^{\top}f_{k_{j}}$
$R(V,X)_{f}=R\left(S,K\right)_{f}=\frac{1}{|V|}\sum_{v_{i}\in S}\max_{k_{j}\in K}f_{v_{i}}^{\top}f_{k_{j}}$
@@
$\text{EVQAScore}(V,X)_{f}=2\frac{P(V,X)_{f}\cdot R(V,X)_{f}}{P(V,X)_{f}+R(V,X)_{f}}$
@@ 4.1 Experimental Setting
We employed the positive-augmented CLIP visual encoder [33] to extract embeddings from individual frames, applying average pooling to represent the entire video. Similarly, we used the corresponding textual encoder to generate token embeddings, producing a comprehensive caption representation. Additionally, we utilized the LLaMA3.1-8B-Instruct model to identify key terms from question-answer pairs, highlighting the essential elements of the dialogue.
@@
We used EMScore, PAC-S, and PAC-S++ as baselines. While all three employ the same computational formula, they differ in the specific CLIP models utilized for their calculations.
@@
For LLaMA3.1-8B-Instruct, we applied the official hyperparameters from the repository. For the CLIP model, we used the clip-ViT-B-32 variant for keyframe selection. All experiments were conducted on a machine equipped with 8 NVIDIA H800 GPUs, a 192-core CPU, and 960GB of memory.
@@ 4.4 Video QA Data Filtering
Table 7: Performance of VideoLLaVA and VideoLLaMA2 on MVBench using filtered data by PAC-S and EVQAScore.
@@ 4.6 Case Study: Evaluating Video Captions
We summarize the first case in Figure 1. Given the question: ”What is the man doing in the video?” the first answer contains only one word and is of low quality, while the second answer is longer and of higher quality. Compared with CLIPScore (using the pooling feature of video), EMScore, PAC-S, and EVQAScore, the video QA pair indicates that our method is the only approach capable of recognizing the second answer as being of higher quality. This demonstrates the robustness of our method in evaluating Video QA data.